a container is filled with neon.the neon escapes through a porous at the rate of 1mole/sec. if the cylinder was emptied and 
filled  25% of hydrogen (by moles), 25% of nitrogen(by moles) and 50%(by moles) of helium what is the rate of effusion of the mixture

Grade:11

1 Answers

Sunil Kumar FP
9 years ago
Here rate of effusion is inversly proportional to its molecular weight.
R=sqrt(P/M)
where p is the partial pressure and Mis the mol.wt
Here effective molecular weigt=(1/2*mol He+1/4*molN2+1/4*mol wt H2)/1
=1/2*4 +1/4*28 +1/4*2=9.5 gm
at.wt of neon=10
now let the rate of effusion be x
thus
x/1=sqrt(20/9.5)
=1.45 mol/sec